What does Mackerel scales and mares' tails make lofty ship carry low sails mean?
It is about cloud formations that predict weather. When the clouds look like a Mackerel's scales with horse tails it predicts high winds which would make it necessary for boats to reef their sails. Hope that helps a bit
Drawing one red ball out of a bag that contains 3 red balls and 4 green balls is 3 in 7.
If we are to randomly select one ball out of the bag and replace it (I assume with the same ball) then there are still 3 red balls and 4 green balls. The odds are still 3 in 7. It doesn't matter how many times you draw. Every time you draw a ball you have a 3 in 7 chance of pulling out a red ball.
The advantages of One-tailed and two-tailed test?
The choice of one-tailed or two-tailed tests follows the logic of the hypothesis that is being tested! The one-tailed test, if appropriate, will be more powerful.
What does the 45 percent tornado probability from the spc typically mean?
This qualifies as a high-risk outlook.
Such an outlook generally means that the SPC anticipates a major tornado outbreak with the potential for multiple long-tracked and violent tornadoes.
Since the SPC started publishing outlooks online in 2003, only four days have featured 45% or greater tornado outlooks.
What are the two major reasons why psychology researchers do not involve probability sample?
Psychology researches do use probability samples, so the question is based on false premises.
All odds prior to a sporting event for betting purposed is based, at least in part on statistical probability. Betting on a card game in different stages, like in Texas Hold 'em, by better players is done on probabilities. Offering a warrantee (guarantee) on manufactured items, like a car. The probability your new car will last 3 years, or 5 years or whatever the warrantee is, is again based on statistical probability. After all they really don't want to pay to fix your car.
Risk analysis. Ex: The manufacturer knows a certain drug will kill a certain number of people. Is it worth the risk of law suits to sell the drug? In other words, will the company make enough profit off of sales to cover the law suits. All Insurance: the probability that a customer will ever cash in on an insurance policy.
What is the probability of rolling a 5 on a roll of a single number cube?
The probability of rolling a 5 on a roll of a single number cube is 1/6 (0.1666) as there is only one number 5 on a total of 6 sides on the cube.
yellow
What is the probability of rolling a number more than 7?
There are 36 permutations of two dice. Of these, 15 add up to a sum greater than 7. The permutations are listed below, with the desired results bolded.
1 + 1 = 2
1 + 2 = 3
1 + 3 = 4
1 + 4 = 5
1 + 5 = 6
1 + 6 = 7
2 + 1 = 3
2 + 2 = 4
2 + 3 = 5
2 + 4 = 6
2 + 5 = 7
2 + 6 = 8
3 + 1 = 4
3 + 2 = 5
3 + 3 = 6
3 + 4 = 7
3 + 5 = 8
3 + 6 = 9
4 + 1 = 5
4 + 2 = 6
4 + 3 = 7
4 + 4 = 8
4 + 5 = 9
4 + 6 = 10
5 + 1 = 6
5 + 2 = 7
5 + 3 = 8
5 + 4 = 9
5 + 5 = 10
5 + 6 = 11
6 + 1 = 7
6 + 2 = 8
6 + 3 = 9
6 + 4 = 10
6 + 5 = 11
6 + 6 = 12
The probability, then, of rolling a sum greater than 7 with two dice is 15 in 36, or about 0.4167.
If six-sided dice are not what you are talking about, then please restate the question and be more specific about what you want to know.
What is empirical probability?
Empirical and experimental probability are the same thing. They are the outcome of an experiment.
Five or less fire calls is: P(0)+P(1)+......+P(5). Formula is: P(x) = (mxe-m)/x! where m=mean & x is number of calls So, P(o) = (3.40e-3.4)/0! = .0334 Likewise P(1)=.1135; P(2)=.1929; P(3)=.2186; P(4)=.1858 & P(5)=.1264. So, P(Five or less fire calls) = .8706
This is the well known Monty Hall problem, named after the presenter of a TV game show which presented players with this dilemma. For a good explanation of the background and winning strategy, see the Wikipedia: Monty Hall problem at: http://en.wikipedia.org/wiki/Monty_Hall_problem
3 red balls and 2 blue balls are in a bag what is the probability that both chosen will be red?
Assuming that you're only drawing two balls out of the bag, your chance of getting two reds would be 3/5 * 2/4 = 3/10, or 30%
What is the probability that the number on the red die is smaller than the number on the white die?
If both cubes are standard 6-sided cubes, then there are six different situations,
depending on what number the white die lands on.
White . . Probability
1 . . . . . . . . 0%
2 . . . . . . . . 16.67%
3 . . . . . . . . 33.33%
4 . . . . . . . . 50%
5 . . . . . . . . 66.67%
6 . . . . . . . . 83.33%
What is independent and dependent when dealing with probability?
The event whose occurrence is not relying on other the other event is independent e.g the occurance of Head in a coin throw is not dependent on other side, the Tail, so it is an independent event.
When two events are depending on each other in order to gain a required result, the events are said to be dependant.
How many times would you chose two fives one after the other in a deck of cards?
The question is a little ambiguous.
The simpler version is: if you picked up a shuffled deck of cards and drew two cards from it, what is the possibility that they would both be fives?
The probability of drawing the first five is 4 out of 52, or 1/13.
The probability of drawing the second five is then 3 out of 51, or 1/17.
Combining these, you get that this should happen on average once in 13x17 = 221 tries.
The more complicated version is: what is the probability that if you KEEP drawing cards from the deck that you will EVENTUALLY pull out two fives in a row? This is somewhat more likely, but the math does get complicated.
How can you use statistical measures to compare populations?
There are different methods for comparing the mean, variance or standard error, distribution or other characteristics of populations. Without more specific information it is not possible to answer the question.
It is 720.
What is the probability of spinning blue on both spinners?
There is insufficient information for us to even begin to answer this question. Please edit the question to include more context or relevant information. There is no information on the shape of the two spinners and what colours are on each.
Yes, to approximately standard normal.
If the random variable X is approximately normal with mean m and standard deviation s, then
(X - m)/s
is approximately standard normal.
What does it mean when something is disjoint or overlapping in math?
Two sets are disjoint if there are elements that belong to both.
Two sets are overlapping if there is at least one elements that belongs to both.